OpenStudy (anonymous):

find the exact value sin(5pi/6). If you could show the steps it would be greatly appreciated.

OpenStudy (anonymous):

step one plug into calculator step two click enter step three you have answer

OpenStudy (anonymous):

step one look at that unit circle on the cheat sheet i sent you step two find \(\frac{5\pi}{6}\) on the circle, then find the coordinates on the circle step three the second coordinate is sine, that is your answer i am not being silly that is how you do this

OpenStudy (anonymous):

Or you can plug it in the calculator? I mean if you're just going to look at a unit circle and figure it out that way might as well use a calculator.

OpenStudy (anonymous):

Thank you, I appreciate it. I went to close the last one because my problem was answered, but after I did that it would not allow me to open it. I did download the Sheet, thank you.

OpenStudy (anonymous):

Ryan, your response is appreciated as well. However, I was trying to grasp a better understanding of how to do it majorly without the calculator.

OpenStudy (anonymous):

5pi/6=(pi-pi/6) sin(pi-x)=sinx therefore sin(5pi/6)=sin(pi/6)=1/2

OpenStudy (anonymous):

unless you have a fancy calculator it will not help your calculator will not display \(\frac{\sqrt{3}}{2}\) for example

OpenStudy (anonymous):

here is the cheat sheet again

OpenStudy (anonymous):

it is on the site "Paul's notes" which is very helpful and has lots of nice cheat sheets as well

OpenStudy (anonymous):

There is always wolfram!!

OpenStudy (anonymous):

yeah, that will work, as long as you don't have to a) learn anything b) take an in class exam

OpenStudy (anonymous):

and wolfram explains everything to you

OpenStudy (anonymous):

\[\sin \frac{ 5\pi }{6 }=\sin \left( \pi-\frac{ \pi }{ 6 } \right)=\sin \frac{ \pi }{ 6 }=\frac{ 1 }{2 }\]

OpenStudy (anonymous):